calculate the zero point energy of system consisting of mass of 10^ -3 kg connected to fixed point by a spring which is stretched by 10^ -2 m by a force of 10 ^ -1 N .
SYSTEM IS CONSTRAINED TO MOVE IN 1 DIRECTION.
1
Expert's answer
2020-01-09T14:03:40-0500

The zero-point energy of system is given by formula

"E=0.5h\u03c9 (1),"

where ω is the oscillator frequency, h=1.055×10-34 J×s


We can calculate the oscillator frequency as

"ω=\sqrt{\frac{k}{m}}= \sqrt{\frac{F}{mx}} (2)"

We put (2) in (1)

"E=0.5h×(\sqrt{\frac{F}{mx}}) (3)"

In our case, we have m=10-3 kg, x=10-2 m, F=10-1 N.

Using (3) we get

E=5.28×10-33 J


Answer

5.28×10-33 J